OpenStudy (anonymous):

A copy machine makes 60 copies per minute. A second copy machine makes 80 copies per minute. The second machine starts making copies 2 minutes after the first machine starts. Both machines stop making copies 8 minutes after the first machine started. Together, the 2 machines made how many copies?

OpenStudy (anonymous):

Hello! I can try to help! Okay, so the first copyer *appologies for crappy spelling long day mate* makes 60 copy's every minute. This can be represented by 60X80=x The second one will be 80X8-160 since it takes two minutes. Add those up and you will have your answer. Cheers!

OpenStudy (anonymous):

Im sure Heybhai will be correct and explain it better though!

OpenStudy (anonymous):

In 1st 2 minutes,machine A produced 60*2=120 copies afterwards both started working together hence for 8 minutes they produced (60+80)*8=140*8=1120 copies Hence total of 120+1120 =1240 copies were made. he point here to understand is as 2 machines started working together their individual performance remains independent of each others work

OpenStudy (anonymous):

how did you get the 6 minutes?